A ball is thrown vertically upwards with an initial speed of 15ms^-1.
i. How long does it take the ball to return to the thrower
ii. What is the maximum height reached by the ball
1
Expert's answer
2019-05-23T09:43:38-0400

With such initial speed the ball will move upward for


"t=v_0\/g."

Then the ball will start falling from the same height and it will achieve the same speed, i.e. this will take the same time and the ball will return to the thrower in


"T=2t=2v_0/g=2\cdot 15/9.8=3.06\text{ s}."

The maximum height will be


"h=v^2/2g=15^2/(2\cdot9.8)=11.48\text{ m}."